exact value
sin(cos^-1 3/8)

Answer by oscargut(686)

sin(cos^-1(x))=sqrt(1-x^2)
then
sin(cos^-1(3/8))=sqrt(1-9/64)
=sqrt(55/64)=(0.927 aprox)
Exact value is sqrt(55/64)